### v1

@@ 6. Conclusion
Time series forecasting is crucial in various fields such as economics, traffic, and AIOps. In this study, we propose a general framework DAG, which utilizes Dual cAusal network along both the temporal and channel dimensions for time series forecasting with exoGenous variables, especially leveraging future exogenous variable information. The framework introduces a Temporal Causal Module which includes a temporal causal discovery module to model how historical exogenous variables affect future exogenous variables, followed by a causal injection module to incorporate these relationships for forecasting future endogenous variables. Additionally, a Channel Causal Module is introduced. It features a channel causal discovery module to model the impact of historical exogenous variables on historical endogenous variables and injects these relationships for improved forecasting of future endogenous variables based on future exogenous variables. Finally, this framework combines temporal causality loss, channel causality loss, and forecasting loss to enable end-to-end optimization. Extensive experiments show that DAG achieves state-of-the-art performance. Additionally, all datasets and code are avaliable at https://github.com/decisionintelligence/DAG.
@@ Appendix A Experimental Details
Table 5. Statistics of datasets. Ex. and En. are abbreviations for the Exogenous variable and Endogenous variable, respectively.
@@
Table 6. Full results on multivariate datasets with TSF-X task, where the inputs are ($X^{\text{endo}},X^{\text{exo}}$, and $Y^{\text{exo}}$). The forecasting horizons $F\in\{96,192,336,720\}$ and look-back window is set to 96. Avg means the average results from all four forecasting horizons. Red: the best, Blue: the 2nd best.
@@
Models DAG (ours) TimeXer TFT TiDE DUET CrossLinear Amplifier TimeKAN xPatch PatchTST DLinear
Metrics mse mae mse mae mse mae mse mae mse mae mse mae mse mae mse mae mse mae mse mae mse mae
ETTh1
96 0.054 0.176 0.056 0.179 0.081 0.214 0.056 0.179 0.163 0.338 0.152 0.321 0.160 0.332 0.118 0.273 0.136 0.305 0.216 0.403 0.176 0.351
192 0.072 0.208 0.073 0.207 0.073 0.213 0.075 0.210 0.206 0.382 0.199 0.374 0.134 0.290 0.166 0.330 0.150 0.311 0.174 0.336 0.227 0.399
336 0.079 0.219 0.089 0.234 0.099 0.246 0.090 0.236 0.248 0.419 0.218 0.383 0.162 0.325 0.140 0.295 0.133 0.291 0.253 0.429 0.187 0.355
720 0.079 0.221 0.104 0.254 0.124 0.282 0.096 0.244 0.283 0.446 0.295 0.475 0.263 0.429 0.284 0.440 0.239 0.406 0.302 0.462 0.350 0.511
Avg 0.071 0.206 0.081 0.219 0.094 0.239 0.079 0.217 0.225 0.396 0.216 0.388 0.180 0.344 0.177 0.335 0.165 0.328 0.236 0.407 0.235 0.404
ETTh2
96 0.127 0.273 0.141 0.289 0.217 0.364 0.127 0.271 0.121 0.275 0.135 0.289 0.119 0.275 0.122 0.278 0.191 0.352 0.134 0.293 0.189 0.343
192 0.177 0.327 0.194 0.346 0.256 0.405 0.179 0.327 0.234 0.391 0.154 0.323 0.200 0.356 0.186 0.354 0.142 0.305 0.158 0.326 0.142 0.303
336 0.215 0.369 0.226 0.380 0.287 0.435 0.223 0.375 0.228 0.379 0.179 0.343 0.209 0.367 0.181 0.342 0.185 0.349 0.172 0.338 0.180 0.339
720 0.214 0.370 0.234 0.385 0.283 0.431 0.253 0.405 0.209 0.365 0.319 0.463 0.348 0.488 0.337 0.473 0.223 0.385 0.337 0.475 0.257 0.405
Avg 0.183 0.335 0.199 0.350 0.261 0.409 0.195 0.344 0.198 0.352 0.197 0.354 0.219 0.372 0.207 0.362 0.185 0.348 0.200 0.358 0.192 0.348
ETTm1
96 0.029 0.127 0.030 0.129 0.030 0.131 0.030 0.130 0.063 0.204 0.087 0.247 0.092 0.252 0.067 0.205 0.055 0.188 0.092 0.253 0.059 0.188
192 0.044 0.159 0.044 0.161 0.048 0.168 0.049 0.168 0.217 0.413 0.164 0.351 0.177 0.359 0.156 0.331 0.128 0.300 0.132 0.304 0.189 0.378
336 0.057 0.184 0.060 0.189 0.068 0.200 0.069 0.203 0.252 0.446 0.239 0.426 0.247 0.435 0.191 0.363 0.155 0.333 0.255 0.440 0.220 0.371
720 0.079 0.216 0.082 0.222 0.095 0.240 0.073 0.211 0.265 0.436 0.286 0.462 0.252 0.424 0.261 0.428 0.247 0.425 0.251 0.434 0.282 0.458
Avg 0.052 0.172 0.054 0.175 0.060 0.185 0.055 0.178 0.199 0.375 0.194 0.372 0.192 0.367 0.169 0.332 0.146 0.312 0.183 0.358 0.187 0.349
ETTm2
96 0.063 0.183 0.072 0.195 0.117 0.253 0.070 0.190 0.107 0.260 0.156 0.332 0.175 0.353 0.146 0.309 0.151 0.320 0.173 0.351 0.113 0.264
192 0.098 0.230 0.110 0.249 0.179 0.324 0.101 0.235 0.200 0.366 0.203 0.366 0.179 0.339 0.139 0.293 0.096 0.235 0.142 0.295 0.156 0.307
336 0.127 0.270 0.139 0.284 0.206 0.356 0.128 0.272 0.136 0.292 0.136 0.292 0.139 0.294 0.182 0.338 0.141 0.291 0.151 0.311 0.137 0.293
720 0.179 0.328 0.190 0.340 0.271 0.417 0.179 0.328 0.147 0.310 0.156 0.316 0.175 0.332 0.169 0.332 0.204 0.362 0.164 0.326 0.152 0.310
Avg 0.117 0.253 0.128 0.267 0.193 0.338 0.119 0.256 0.148 0.307 0.163 0.326 0.167 0.330 0.159 0.318 0.148 0.302 0.157 0.321 0.139 0.294
Weather
96 0.001 0.023 0.001 0.023 0.001 0.028 0.001 0.026 0.002 0.028 0.012 0.093 0.008 0.074 0.008 0.072 0.001 0.027 0.006 0.062 0.008 0.076
192 0.001 0.022 0.001 0.026 0.002 0.030 0.001 0.028 0.002 0.030 0.005 0.056 0.008 0.073 0.009 0.078 0.001 0.024 0.008 0.073 0.006 0.065
336 0.002 0.028 0.001 0.027 0.002 0.030 0.002 0.031 0.001 0.025 0.009 0.080 0.008 0.071 0.009 0.074 0.002 0.029 0.011 0.087 0.010 0.081
720 0.002 0.030 0.002 0.032 0.002 0.035 0.002 0.036 0.002 0.033 0.007 0.067 0.019 0.113 0.003 0.045 0.002 0.037 0.012 0.084 0.014 0.098
Avg 0.002 0.026 0.001 0.027 0.002 0.031 0.002 0.030 0.002 0.029 0.008 0.074 0.011 0.083 0.007 0.067 0.002 0.029 0.009 0.076 0.010 0.080
Electricity
96 0.158 0.298 0.160 0.296 0.365 0.456 0.312 0.414 0.248 0.385 0.281 0.416 0.279 0.416 0.276 0.413 0.249 0.387 0.274 0.406 0.301 0.433
192 0.156 0.295 0.186 0.320 0.339 0.448 0.274 0.391 0.290 0.421 0.248 0.384 0.296 0.427 0.309 0.435 0.258 0.397 0.264 0.399 0.303 0.427
336 0.217 0.353 0.218 0.346 0.367 0.471 0.313 0.417 0.295 0.426 0.260 0.396 0.278 0.414 0.292 0.419 0.296 0.425 0.271 0.408 0.253 0.396
720 0.246 0.370 0.267 0.386 0.429 0.510 0.353 0.445 0.290 0.414 0.268 0.401 0.255 0.396 0.327 0.448 0.256 0.392 0.268 0.402 0.296 0.422
Avg 0.194 0.329 0.208 0.337 0.375 0.471 0.313 0.417 0.281 0.411 0.264 0.399 0.277 0.413 0.301 0.429 0.265 0.400 0.269 0.404 0.288 0.420
Traffic
96 0.119 0.187 0.135 0.233 0.139 0.221 0.149 0.251 0.247 0.316 0.261 0.322 0.174 0.240 0.221 0.292 0.244 0.305 0.241 0.307 0.185 0.258
192 0.121 0.191 0.135 0.230 0.116 0.212 0.149 0.246 0.240 0.297 0.232 0.289 0.214 0.276 0.181 0.254 0.220 0.292 0.246 0.305 0.240 0.306
336 0.125 0.200 0.143 0.237 0.134 0.228 0.150 0.251 0.310 0.356 0.223 0.280 0.188 0.318 0.210 0.283 0.211 0.278 0.215 0.280 0.241 0.313
720 0.152 0.232 0.156 0.250 0.138 0.220 0.169 0.271 0.253 0.318 0.213 0.272 0.201 0.271 0.231 0.310 0.210 0.273 0.229 0.286 0.183 0.247
Avg 0.129 0.203 0.142 0.238 0.132 0.220 0.154 0.255 0.263 0.322 0.232 0.291 0.194 0.276 0.211 0.285 0.221 0.287 0.233 0.295 0.212 0.281
Exchange
96 0.087 0.229 0.109 0.245 0.105 0.258 0.096 0.230 0.291 0.484 0.234 0.425 0.136 0.316 0.297 0.487 0.220 0.397 0.370 0.554 0.113 0.283
192 0.147 0.298 0.197 0.350 0.258 0.393 0.205 0.340 0.236 0.421 0.231 0.430 0.186 0.360 0.300 0.461 0.153 0.334 0.427 0.590 0.363 0.525
336 0.274 0.413 0.331 0.444 0.422 0.524 0.453 0.519 0.309 0.496 0.394 0.563 0.412 0.562 0.716 0.752 0.484 0.610 0.558 0.687 0.427 0.576
720 0.338 0.470 0.362 0.482 0.552 0.577 0.419 0.502 0.244 0.419 0.390 0.558 0.459 0.614 0.387 0.559 0.336 0.511 0.503 0.640 0.550 0.655
Avg 0.212 0.353 0.250 0.380 0.334 0.438 0.293 0.398 0.270 0.455 0.312 0.494 0.298 0.463 0.425 0.565 0.298 0.463 0.464 0.618 0.363 0.510
$1^{st}$ Count 28 28 5 4 2 1 1 2 5 4 0 0 1 0 0 0 2 0 1 1 0 1
A.1. Datasets

### v2

@@ 6 Conclusion
In this study, we propose a general framework DAG, which utilizes Dual correlAtion network along both the temporal and channel dimensions for time series forecasting with exoGenous variables, especially leveraging future exogenous variable information. The framework introduces a Temporal Correlation Module which includes a temporal correlation discovery module to model how historical exogenous variables affect future exogenous variables, followed by a correlation injection module to incorporate these relationships for forecasting future endogenous variables. Additionally, a Channel Correlation Module is introduced. It features a channel correlation discovery module to model the impact of historical exogenous variables on historical endogenous variables and injects these relationships for improved forecasting of future endogenous variables based on future exogenous variables. Finally, this framework combines temporal correlation loss, channel correlation loss, and forecasting loss to enable end-to-end optimization. Additionally, all datasets and code are available at https://github.com/decisionintelligence/DAG.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Experimental Details
Table 4: Statistics of datasets. Ex. and En. are abbreviations for the Exogenous variable and Endogenous variable, respectively.
@@
Table 5: Results on 12 real-world datasets that satisfy the TSF-X task, where the inputs are ($X^{\text{endo}},X^{\text{exo}}$, and $Y^{\text{exo}}$). Red: the best, Blue: the 2nd best. Avg means the average results from two forecasting horizons.
@@
Models DAG (ours) TimeXer TFT TiDE DUET CrossLinear Amplifier TimeKAN PatchTST DLinear
Metrics mse mae mse mae mse mae mse mae mse mae mse mae mse mae mse mae mse mae mse mae
NP 24 0.202 0.237 0.236 0.266 0.219 0.249 0.284 0.301 0.246 0.287 0.210 0.266 0.252 0.303 0.273 0.310 0.249 0.294 0.282 0.337
360 0.521 0.451 0.600 0.475 0.539 0.501 0.601 0.498 0.576 0.528 0.531 0.508 0.587 0.534 0.538 0.529 0.530 0.498 0.601 0.552
Avg 0.362 0.344 0.418 0.371 0.379 0.375 0.443 0.400 0.411 0.408 0.371 0.387 0.420 0.418 0.405 0.419 0.390 0.396 0.442 0.444
PJM 24 0.057 0.143 0.075 0.166 0.095 0.195 0.106 0.214 0.072 0.166 0.088 0.191 0.096 0.208 0.115 0.244 0.116 0.239 0.096 0.208
360 0.130 0.218 0.140 0.231 0.133 0.219 0.177 0.279 0.131 0.228 0.135 0.254 0.177 0.285 0.162 0.281 0.149 0.279 0.164 0.278
Avg 0.093 0.180 0.108 0.198 0.114 0.207 0.142 0.246 0.102 0.197 0.112 0.223 0.137 0.246 0.139 0.262 0.133 0.259 0.130 0.243
BE 24 0.361 0.229 0.392 0.253 0.426 0.272 0.426 0.285 0.432 0.272 0.391 0.259 0.471 0.339 0.451 0.319 0.452 0.326 0.451 0.310
360 0.485 0.330 0.512 0.327 0.482 0.310 0.571 0.364 0.597 0.436 0.568 0.416 0.646 0.487 0.645 0.495 0.702 0.538 0.600 0.442
Avg 0.423 0.279 0.452 0.290 0.454 0.291 0.498 0.325 0.515 0.354 0.479 0.337 0.559 0.413 0.548 0.407 0.577 0.432 0.526 0.376
FR 24 0.355 0.171 0.366 0.208 0.543 0.253 0.418 0.255 0.384 0.251 0.390 0.226 0.459 0.348 0.454 0.296 0.518 0.368 0.422 0.265
360 0.473 0.268 0.489 0.273 0.465 0.261 0.551 0.308 0.607 0.403 0.575 0.370 0.648 0.468 0.641 0.452 0.658 0.452 0.572 0.362
Avg 0.414 0.219 0.427 0.241 0.504 0.257 0.484 0.281 0.496 0.327 0.483 0.298 0.554 0.408 0.547 0.374 0.588 0.410 0.497 0.314
DE 24 0.277 0.322 0.339 0.362 0.380 0.383 0.367 0.383 0.376 0.378 0.387 0.396 0.394 0.407 0.399 0.412 0.413 0.412 0.423 0.426
360 0.462 0.418 0.610 0.474 0.599 0.509 0.630 0.511 0.589 0.482 0.583 0.507 0.551 0.474 0.547 0.479 0.589 0.498 0.573 0.487
Avg 0.370 0.370 0.475 0.418 0.489 0.446 0.499 0.447 0.482 0.430 0.485 0.452 0.473 0.441 0.473 0.445 0.501 0.455 0.498 0.457
Energy 24 0.079 0.215 0.122 0.273 0.093 0.235 0.103 0.248 0.117 0.283 0.241 0.418 0.138 0.306 0.135 0.298 0.239 0.390 0.085 0.230
360 0.169 0.320 0.204 0.357 0.167 0.331 0.202 0.355 0.288 0.452 0.237 0.385 0.328 0.472 0.302 0.464 0.214 0.363 0.263 0.416
Avg 0.124 0.267 0.163 0.315 0.130 0.283 0.153 0.302 0.203 0.367 0.239 0.402 0.233 0.389 0.218 0.381 0.226 0.377 0.174 0.323
Sdwpfm1 24 0.351 0.400 0.558 0.533 0.366 0.421 0.474 0.488 0.551 0.564 0.355 0.473 0.364 0.445 0.418 0.503 0.374 0.454 0.417 0.501
360 0.495 0.522 0.845 0.684 0.597 0.528 0.492 0.526 0.646 0.577 0.497 0.532 0.510 0.535 0.476 0.566 0.497 0.550 0.480 0.562
Avg 0.423 0.461 0.701 0.609 0.482 0.474 0.483 0.507 0.599 0.570 0.426 0.502 0.437 0.490 0.447 0.534 0.435 0.502 0.449 0.531
Sdwpfm2 24 0.372 0.414 0.627 0.570 0.411 0.458 0.461 0.492 0.445 0.452 0.477 0.536 0.394 0.462 0.474 0.538 0.418 0.492 0.459 0.530
360 0.583 0.556 0.978 0.736 0.541 0.519 0.511 0.540 0.584 0.528 0.589 0.611 0.587 0.563 0.520 0.589 0.602 0.603 0.589 0.630
Avg 0.477 0.485 0.803 0.653 0.476 0.488 0.486 0.516 0.514 0.490 0.533 0.573 0.491 0.512 0.497 0.564 0.510 0.547 0.524 0.580
Sdwpfh1 24 0.408 0.438 0.651 0.587 0.401 0.460 0.434 0.489 0.527 0.513 0.548 0.585 0.576 0.627 0.511 0.582 0.422 0.505 0.557 0.613
360 0.489 0.534 0.841 0.700 0.557 0.523 0.472 0.527 0.551 0.519 0.566 0.601 0.497 0.569 0.643 0.694 0.513 0.548 0.600 0.635
Avg 0.448 0.486 0.746 0.643 0.479 0.491 0.453 0.508 0.539 0.516 0.557 0.593 0.537 0.598 0.577 0.638 0.468 0.527 0.579 0.624
Sdwpfh2 24 0.438 0.465 0.820 0.677 0.474 0.493 0.579 0.553 0.629 0.563 0.468 0.540 0.473 0.533 0.580 0.614 0.457 0.527 0.633 0.653
360 0.608 0.595 0.962 0.761 0.657 0.549 0.619 0.614 0.665 0.569 0.608 0.609 0.569 0.628 0.713 0.729 0.641 0.632 0.704 0.674
Avg 0.523 0.530 0.891 0.719 0.566 0.521 0.599 0.583 0.647 0.566 0.538 0.574 0.521 0.581 0.647 0.672 0.549 0.580 0.669 0.664
Colbun 10 0.061 0.094 0.113 0.172 0.092 0.135 0.089 0.131 0.089 0.134 0.071 0.102 0.071 0.121 0.061 0.101 0.062 0.122 0.069 0.151
30 0.135 0.215 0.176 0.299 0.383 0.460 0.240 0.322 0.307 0.397 0.182 0.288 0.275 0.370 0.195 0.249 0.417 0.496 0.245 0.285
Avg 0.098 0.154 0.145 0.235 0.238 0.297 0.164 0.227 0.198 0.266 0.126 0.195 0.173 0.246 0.128 0.175 0.239 0.309 0.157 0.218
Rapel 10 0.151 0.203 0.301 0.308 0.201 0.253 0.228 0.271 0.174 0.219 0.163 0.209 0.181 0.227 0.174 0.231 0.163 0.218 0.184 0.239
30 0.309 0.408 0.387 0.416 0.409 0.414 0.411 0.432 0.365 0.432 0.340 0.417 0.333 0.416 0.325 0.390 0.374 0.445 0.316 0.410
Avg 0.230 0.305 0.344 0.362 0.305 0.333 0.320 0.351 0.269 0.326 0.252 0.313 0.257 0.321 0.249 0.311 0.269 0.332 0.250 0.324
1st Count 26 29 0 0 5 5 2 0 0 1 0 0 2 0 1 1 0 0 0 0
Table 6: Results under the setting where future exogenous variables are not available. The inputs are ($X^{\text{endo}}$ and $X^{\text{exo}}$). The best results are Red, and the second-best restuls are Blue. Avg represents the average results across the two forecasting horizons.
Models DAG TimeXer TFT TiDE DUET CrossLinear Amplifier TimeKAN PatchTST DLinear
Metrics mse mae mse mae mse mae mse mae mse mae mse mae mse mae mse mae mse mae mse mae
NP 24 0.237 0.259 0.270 0.292 0.330 0.321 0.297 0.309 0.262 0.271 0.246 0.282 0.302 0.321 0.336 0.344 0.283 0.296 0.294 0.308
360 0.601 0.500 0.609 0.474 0.964 0.655 0.638 0.523 0.626 0.494 0.655 0.505 0.738 0.574 0.632 0.527 0.630 0.505 0.647 0.537
Avg 0.419 0.380 0.440 0.383 0.647 0.488 0.467 0.416 0.444 0.383 0.451 0.394 0.520 0.448 0.484 0.435 0.457 0.401 0.471 0.423
PJM 24 0.072 0.162 0.096 0.191 0.129 0.231 0.105 0.213 0.082 0.178 0.096 0.198 0.093 0.195 0.130 0.236 0.106 0.212 0.109 0.215
360 0.180 0.259 0.186 0.267 0.270 0.310 0.210 0.292 0.198 0.273 0.197 0.283 0.210 0.294 0.219 0.304 0.189 0.273 0.199 0.285
Avg 0.126 0.210 0.141 0.229 0.200 0.270 0.158 0.253 0.140 0.226 0.147 0.241 0.152 0.245 0.175 0.270 0.148 0.243 0.154 0.250
BE 24 0.376 0.244 0.392 0.253 0.400 0.267 0.492 0.322 0.376 0.253 0.389 0.253 0.417 0.283 0.424 0.280 0.405 0.264 0.453 0.308
360 0.547 0.350 0.563 0.349 0.727 0.442 0.602 0.375 0.571 0.357 0.564 0.346 0.588 0.367 0.552 0.350 0.564 0.354 0.645 0.454
Avg 0.462 0.297 0.477 0.301 0.563 0.354 0.547 0.348 0.473 0.305 0.477 0.300 0.502 0.325 0.488 0.315 0.485 0.309 0.549 0.381
FR 24 0.345 0.190 0.366 0.195 0.445 0.231 0.415 0.254 0.359 0.208 0.397 0.204 0.429 0.250 0.448 0.250 0.397 0.223 0.412 0.253
360 0.526 0.307 0.542 0.300 0.624 0.345 0.574 0.325 0.577 0.317 0.555 0.310 0.559 0.322 0.534 0.303 0.542 0.306 0.576 0.352
Avg 0.435 0.249 0.454 0.247 0.535 0.288 0.494 0.290 0.468 0.262 0.476 0.257 0.494 0.286 0.491 0.276 0.470 0.264 0.494 0.302
DE 24 0.422 0.411 0.501 0.445 0.576 0.452 0.465 0.433 0.424 0.403 0.438 0.423 0.493 0.455 0.504 0.462 0.503 0.450 0.511 0.458
360 0.784 0.563 0.818 0.568 0.792 0.578 0.823 0.604 0.895 0.623 0.832 0.592 0.930 0.642 0.834 0.590 0.889 0.604 0.868 0.600
Avg 0.603 0.487 0.659 0.507 0.684 0.515 0.644 0.519 0.660 0.513 0.635 0.508 0.712 0.548 0.669 0.526 0.696 0.527 0.690 0.529
Energy 24 0.112 0.257 0.138 0.293 0.352 0.463 0.117 0.265 0.108 0.254 0.102 0.246 0.108 0.254 0.109 0.254 0.108 0.254 0.106 0.248
360 0.189 0.342 0.206 0.360 0.399 0.497 0.207 0.358 0.211 0.362 0.299 0.425 0.251 0.400 0.186 0.338 0.298 0.428 0.214 0.367
Avg 0.150 0.300 0.172 0.326 0.376 0.480 0.162 0.311 0.160 0.308 0.201 0.336 0.180 0.327 0.147 0.296 0.203 0.341 0.160 0.307
Sdwpfm1 24 0.532 0.512 0.558 0.533 0.596 0.531 0.561 0.540 0.550 0.495 0.549 0.553 0.579 0.557 0.598 0.579 0.557 0.557 0.547 0.589
360 0.845 0.695 0.845 0.684 1.372 0.925 0.864 0.725 0.898 0.672 1.070 0.834 1.106 0.814 0.853 0.764 0.910 0.745 0.839 0.773
Avg 0.689 0.604 0.701 0.609 0.984 0.728 0.713 0.633 0.724 0.583 0.809 0.694 0.843 0.685 0.725 0.672 0.733 0.651 0.693 0.681
Sdwpfm2 24 0.609 0.548 0.627 0.570 0.749 0.657 0.636 0.582 0.612 0.537 0.635 0.586 0.669 0.600 0.684 0.621 0.648 0.610 0.631 0.631
360 0.962 0.766 0.978 0.736 1.340 0.877 0.995 0.783 1.028 0.745 0.965 0.788 1.215 0.863 0.987 0.780 1.020 0.819 0.980 0.824
Avg 0.786 0.657 0.803 0.653 1.044 0.767 0.816 0.682 0.820 0.641 0.800 0.687 0.942 0.732 0.836 0.701 0.834 0.714 0.805 0.728
Sdwpfh1 24 0.647 0.592 0.651 0.587 0.754 0.670 0.719 0.641 0.673 0.575 0.687 0.652 1.111 0.816 0.755 0.679 0.669 0.629 0.736 0.721
360 0.820 0.695 0.841 0.700 0.831 0.727 0.897 0.757 0.886 0.712 0.849 0.752 1.089 0.823 0.853 0.762 0.939 0.804 0.829 0.781
Avg 0.733 0.643 0.746 0.643 0.793 0.698 0.808 0.699 0.779 0.644 0.768 0.702 1.100 0.820 0.804 0.720 0.804 0.717 0.783 0.751
Sdwpfh2 24 0.739 0.635 0.820 0.677 0.811 0.714 0.783 0.674 0.975 0.694 0.810 0.718 0.834 0.691 0.892 0.740 0.769 0.680 0.838 0.759
360 1.000 0.755 0.962 0.761 1.042 0.779 1.056 0.829 1.039 0.735 1.102 0.830 1.126 0.845 0.990 0.825 1.280 0.925 0.989 0.837
Avg 0.870 0.695 0.891 0.719 0.926 0.746 0.919 0.751 1.007 0.715 0.956 0.774 0.980 0.768 0.941 0.782 1.025 0.802 0.913 0.798
Colbun 10 0.072 0.103 0.084 0.146 0.449 0.279 0.095 0.140 0.073 0.106 0.068 0.101 0.077 0.120 0.091 0.129 0.083 0.125 0.077 0.135
30 0.161 0.208 0.181 0.293 0.663 0.494 0.282 0.341 0.221 0.290 0.190 0.289 0.227 0.301 0.311 0.377 0.217 0.316 0.318 0.359
Avg 0.117 0.155 0.132 0.219 0.556 0.386 0.188 0.240 0.147 0.198 0.129 0.195 0.152 0.210 0.201 0.253 0.150 0.221 0.198 0.247
Rapel 10 0.195 0.215 0.301 0.308 0.238 0.274 0.238 0.270 0.221 0.225 0.201 0.224 0.240 0.259 0.226 0.238 0.215 0.232 0.223 0.253
30 0.327 0.347 0.387 0.416 0.377 0.394 0.409 0.444 0.387 0.386 0.278 0.374 0.433 0.436 0.458 0.437 0.436 0.436 0.414 0.423
Avg 0.261 0.281 0.344 0.362 0.308 0.334 0.323 0.357 0.304 0.306 0.240 0.299 0.337 0.348 0.342 0.337 0.325 0.334 0.319 0.338
1st Count 28 19 1 4 0 0 0 0 0 8 4 3 0 0 2 2 0 0 1 0
A.1 Datasets